# Tomographic constraint on anisotropic cosmic birefringence

### v1

@@ Abstract
We for the first time constrain anisotropic cosmic birefringence generated at reionization using Planck PR4 polarization data. Several recent analyses of WMAP and Planck polarization data have found a tantalizing hint of isotropic cosmic birefringence. Ongoing and future CMB experiments will test isotropic cosmic birefringence by improving the absolute angle calibration and understanding the intrinsic parity-odd power spectrum of the Galactic foregrounds. Alternatively, measuring anisotropies in cosmic birefringence and its time evolution is also a key observable to confirm the signal of cosmic birefringence and to investigate its origin. We discuss estimators of anisotropic cosmic birefringence generated at different redshifts. We then estimate anisotropic cosmic birefringence generated at reionization from the PR4 data, showing that the power spectrum is consistent with null. We find that the model proposed by Ferreira et al. (2024) is still consistent with the observation. Future full-sky CMB experiments such as LiteBIRD and PICO will help tighten the tomographic constraint to test models of cosmic birefringence.
@@ I Introduction
where $g_{\phi\gamma}$ is the coupling constant, $\phi$ is an ALP field, $F^{\mu\nu}$ denotes the electromagnetic field tensor, and $\tilde{F}_{\mu\nu}$ is its dual. Multiple studies have explored cosmic birefringence induced by the ALP field of dark energy Carroll:1998:DE ; Liu:2006:biref-time-evolve ; Panda:2010 ; Fujita:2020aqt ; Fujita:2020ecn ; Choi:2021aze ; Obata:2021 ; Gasparotto:2022uqo ; Galaverni:2023 , early dark energy Fujita:2020ecn ; Murai:2022:EDE ; Eskilt:2023:EDE ; Kochappan:2024:biref , dark matter Finelli:2009 ; Sigl:2018:biref-sup ; Liu:2016dcg ; Fedderke:2019:biref ; Zhang:2024dmi , and by topological defects Takahashi:2020tqv ; Kitajima:2022jzz ; Jain:2022jrp ; Gonzalez:2022mcx , as well as by potential signatures of quantum gravity Myers:2003fd ; Balaji:2003sw ; Arvanitaki:2009fg . Looking ahead, future CMB experiments such as BICEP Cornelison:2022:BICEP3 ; BICEPArray , Simons Array SimonsArray , Simons Observatory SimonsObservatory , CMB-S4 CMBS4 , and LiteBIRD LiteBIRD are expected to significantly reduce polarization noise and enhance sensitivity to cosmic birefringence signals.
@@
To confirm these signals, further work is needed to address the intrinsic parity-odd correlations in Galactic foregrounds, which may be used to calibrate the absolute polarization angle Huffenberger:2019:dustEB ; Cukierman:2022:dust ; Hervias-Caimapo:2024 . Another more direct approach is to reduce uncertainties in the absolute polarization angle by improving calibration techniques SO:wiregrid ; Ritacco:2023mgi . Alternatively, observables unaffected by miscalibration angle can be considered. For example, recent studies indicate that the time evolution of pseudoscalar fields during recombination and reionization significantly alters the CMB polarization power spectra Finelli:2009 ; Lee:2013:const ; Gubitosi:2014:biref-time ; Sherwin:2021:biref ; Nakatsuka:2022 ; Naokawa:2023 ; Yin:2023:biref ; Naokawa:2024xhn ; Murai:2024yul . Measuring the spectral shape of the $EB$ power spectrum could break the degeneracy between the miscalibration angle and cosmic birefringence signals. Additional constraints on the late-time evolution of ALPs could be obtained by observing the polarized Sunyaev-Zel’dovich effect Lee:2022:pSZ-biref ; Namikawa:2023:pSZ and galaxy polarization Carroll:1997:radio ; Yin:2024:galaxy . These tomographic signals can help reduce degeneracies with instrumental miscalibration angles.
@@
Anisotropies in cosmic birefringence are another critical observable that can confirm the isotropic signal. If the ALP field exhibits spatial fluctuations, $\delta\phi$, the resulting polarization rotation will be anisotropic (see, e.g., Carroll:1998:DE ; Lue:1999:biref-EB ; Caldwell:2011 ; Lee:2015 ; Leon:2017 ; Yin:2023:biref ; Ferreira:2023 ). Several studies have placed upper limits on these anisotropies by reconstructing the rotation angle Gluscevic:2012 ; PB15:rot ; BICEP2:2017lpa ; Contreras:2017 ; Namikawa:2020:biref ; SPT:2020:biref ; Gruppuso:2020 ; Bortolami:2022whx ; BK-LoS:2023 ; Zagatti:2024 , while others have constrained them using CMB polarization power spectra Li:2014 ; Alighieri:2014yoa ; Liu:2016dcg ; Zhang:2024dmi . However, time evolution tends to significantly suppress the $B$-mode power spectrum significantly Namikawa:2024:BB .
@@
This paper is organized as follows. In Sec. II, we explain our reconstruction methodology. In Sec. III, we describe our data and simulation and test the estimator of anisotropic cosmic birefringence generated during reionization. Sec. IV shows our results for the reconstructed spectrum and the resulting constraint on the scale-invariant birefringence spectrum. We summarize our results and provide a forecast in Sec. V.
@@ II Tomographic reconstruction of anisotropic cosmic birefrigence
For a single source plane, such as the last-scattering surface of CMB, the rotation angle field, $\alpha(\chi_{*},\hat{\bm{n}})$, can be reconstructed from the off-diagonal mode-mode covariance within and between, the $E$ and $B$ modes Kamionkowski:2009:derot ; Namikawa:2020:biref . The power spectrum of the anisotropic rotation angle $C_{L}^{\alpha\alpha}$ can be obtained by squaring the rotation estimator and subtracting relevant biases Namikawa:2016:biref-est . Here, we extend the cases if the CMB photons are generated at multiple epochs.
@@ II.1 Anisotropic cosmic birefringence from recombination and reionization
As a simple case, we first consider if the CMB photons are generated from the recombination and reionization epochs. The presence of the cosmic birefringence effect rotates the primordial Stokes parameters as Greco:2022:aniso-biref-tomography
@@
Consequently, the rotation angle modifies the CMB $E$ and $B$ modes. The $E$ and $B$ modes are obtained by transforming $Q$ and $U$ maps with the spin-2 spherical harmonics, $Y_{\ell m}^{\pm 2}$, as Kamionkowski:1996:eb ; Zaldarriaga:1996xe :
@@
Thus, the $E$ and $B$ modes in the presence of an anisotropic rotation angle are derived by substituting Eq. 2 into the above equation and are given up to linear order in $\alpha$ by:
@@
Here, $p^{\pm}_{\ell_{1}\ell_{2}\ell_{3}}=[1\pm(-1)^{\ell_{1}+\ell_{2}+\ell_{3}}]/2$ is a parity indicator. The ending parentheses denote the Wigner $3j$-symbol. The off-diagonal elements of the covariance induced by the anisotropies of the rotation angle are given by
@@
Here, $\overline{E}_{\ell m}$ and $\overline{B}_{\ell m}$ are the observed multipoles filtered by their inverse variance. For a diagonal filtering, we obtain $\overline{X}_{\ell m}=\widehat{X}_{\ell m}/\widehat{C}^{\rm XX}_{\ell}$, where $X$ is either $E$ or $B$ and $\widehat{C}^{\rm XX}_{\ell}$ is the power spectrum of the observed multipoles, $\widehat{X}_{\ell m}$. Finally, we correct for the mean-field bias, $\langle\bar{\alpha}^{\rm rei}_{LM}\rangle$, and normalize to obtain the rotation angle:
@@
We can compute Eqs. 8 and 10 with a computationally efficient method as described in Appendix A of Namikawa:2020:biref .
@@ II.2 Anisotropic cosmic birefringence from multiple redshifts
Analogues to the estimator for $\alpha^{\rm rei}$, we define the estimator for $\alpha^{z}$ using Eq. (12) as follows:
@@
However, this estimator is biased because its expectation value includes contributions from anisotropic birefringence at different redshifts:
@@ III Data and simulations
Figure 1: Angular power spectrum of anisotropic cosmic birefringence at reionization measured from the Planck PR4 simulation with the $70\%$ Galactic mask. The CMB polarization and noise are derived from the Commander component separation method. The error bars denote the standard deviation with $400$ realizations. We also show the mean-field bias (orange) and disconnected bias, $N^{(0)}$ (green), which corresponds to the noise power spectrum of $\alpha^{\rm rei}_{LM}$.
Figure 2: Correlation coefficients of the angular power spectrum of anisotropic cosmic birefringence measured from the Planck PR4 simulation, as shown in Fig. 1. The lower-left triangle displays the results when using the realization-dependent disconnected bias, while the upper-left triangle shows, for comparison, the case where the simulation average of the disconnected bias is used.
@@
We use the Planck PR4 polarization data to reconstruct anisotropic cosmic birefringence, employing the curved-sky quadratic estimator to extract large-scale birefringence anisotropies.  [^1] The code for reconstructing cosmic birefringence on the full sky is available at https://toshiyan.github.io/clpdoc/html/.
The Planck PR4 polarization map is produced using two different component separation algorithms: Commander and SEVEM. We use both maps and the corresponding simulation sets to reconstruct anisotropic cosmic birefringence. Additionally, we generate a simulation where the CMB signal components are replaced with a signal that includes non-zero anisotropic cosmic birefringence, whose power spectrum follows a scale-invariant model. In this simulation, we first generate $\alpha_{LM}^{\rm rei}$ from a scale-invariant spectrum $C_{L}^{\alpha\alpha}=A_{\rm CB}^{\rm rei}\times 2\pi/L(L+1)$, and then rotate the CMB polarization signals at $\ell\leq 10$ with this rotation fields.
@@
We evaluate the mean-field bias by averaging over the standard simulation realizations. Note that mean-field bias is also induced by the miscalibration angle Namikawa:2020:biref , which is not included in the simulations provided by the Planck collaboration. We test this effect by rotating the polarization map in the simulation as we discuss in the next section.
In practice, partial sky coverage introduces mixing between different multipoles, and the diagonal normalization does not provide an accurate normalization. To correct for this, we follow the approach used in CMB lensing analyses. We first reconstruct $\alpha$ from a simulation that includes non-zero anisotropic birefringence. Next, we cross-correlate the input $\alpha$ with the reconstructed $\alpha$ and rescale the normalization at each $L$ so that the cross-power spectrum agrees with the input birefringence power spectrum.
From the reconstructed $\alpha$, we estimate the cosmic birefringence spectrum following Ref. Namikawa:2016:biref-est . The power spectrum of the estimator defined in Eq. 9 is a four-point function and is decomposed into the disconnected and connected parts. The disconnected part, $N_{L}^{(0)}$, is non-zero even in the absence of birefringence. We evaluate the disconnected bias using the realization-dependent approach Namikawa:2020:biref and subtract it from the power spectrum of the estimator. The N1 is evaluated using simulation from the non-zero birefringence simulations. We also regard the residual power spectrum in the PR4 standard simulation as other biases from the connected four-point function, $N^{\rm MC}_{L}$, arising from lensing Namikawa:2016:biref-est , partial-sky coverage, and noise.
Fig. 1 shows the angular power spectrum of anisotropic cosmic birefringence from the PR4 simulation, which contains no birefringence signals. The estimated power spectrum is consistent with a null result. Fig. 2 plots the correlation coefficients of the angular power spectrum. The realization-dependent method for $N_{L}^{(0)}$ successfully reduces the off-diagonal covariance. Fig. 3 shows the angular power spectrum obtained from a simulation that includes anisotropic cosmic birefringence signals. The power spectrum estimator accurately reproduces the input angular power spectrum. The mean-field, $N^{(1)}$, and $N^{\rm MC}$ are smaller than or comparable to the input signal with $A_{\rm CB}^{\rm rei}=2\times 10^{-3}$ which is much lower than the value constrained from the real data as we show in the next section.

### v2

@@ Abstract
We constrain anisotropic cosmic birefringence generated at reionization using Planck PR4 polarization data for the first time. Several recent analyses of WMAP and Planck polarization data have found a tantalizing hint of isotropic cosmic birefringence. Ongoing and future cosmic microwave background (CMB) experiments will test isotropic cosmic birefringence by improving the absolute angle calibration and understanding the intrinsic parity-odd power spectrum of the Galactic foregrounds. Alternatively, measuring anisotropies in cosmic birefringence and its time evolution is also a key observable to confirm the signal of cosmic birefringence and to investigate its origin. We discuss estimators of anisotropic cosmic birefringence generated at different redshifts. We then estimate anisotropic cosmic birefringence generated at reionization from the PR4 data, showing that the power spectrum is consistent with null. We find that the model proposed by Ferreira et al. is still consistent with the observation. Future full-sky CMB experiments such as LiteBIRD and PICO will help tighten the tomographic constraint to test models of cosmic birefringence.
@@ I Introduction
where $g_{\phi\gamma}$ is the coupling constant, $\phi$ is an ALP field, $F^{\mu\nu}$ denotes the electromagnetic field tensor, and $\tilde{F}_{\mu\nu}$ is its dual. Multiple studies have explored cosmic birefringence induced by the ALP field of dark energy Carroll:1998:DE ; Liu:2006:biref-time-evolve ; Panda:2010 ; Fujita:2020aqt ; Fujita:2020ecn ; Choi:2021aze ; Obata:2021 ; Gasparotto:2022uqo ; Galaverni:2023 , early dark energy Fujita:2020ecn ; Murai:2022:EDE ; Eskilt:2023:EDE ; Kochappan:2024:biref , dark matter Finelli:2009 ; Sigl:2018:biref-sup ; Liu:2016dcg ; Fedderke:2019:biref ; Zhang:2024dmi , and topological defects Takahashi:2020tqv ; Kitajima:2022jzz ; Jain:2022jrp ; Gonzalez:2022mcx , as well as by potential signatures of quantum gravity Myers:2003fd ; Balaji:2003sw ; Arvanitaki:2009fg . Looking ahead, future CMB experiments such as BICEP Cornelison:2022:BICEP3 ; BICEPArray , Simons Array SimonsArray , Simons Observatory SimonsObservatory , CMB-S4 CMBS4 , and LiteBIRD LiteBIRD are expected to significantly reduce polarization noise and enhance sensitivity to cosmic birefringence signals.
@@
To confirm these signals, further work is needed to address the intrinsic parity-odd correlations in Galactic foregrounds, which may be used to calibrate the absolute polarization angle Huffenberger:2019:dustEB ; Cukierman:2022:dust ; Hervias-Caimapo:2024 . Another more direct approach is to reduce uncertainties in the absolute polarization angle by improving calibration techniques SO:wiregrid ; Coppi:2022 ; Ritacco:2023mgi . Alternatively, observables unaffected by miscalibrated polarization angle can be considered. For example, recent studies indicate that the time evolution of pseudoscalar fields during recombination and reionization significantly alters the CMB polarization power spectra Finelli:2009 ; Lee:2013:const ; Gubitosi:2014:biref-time ; Sherwin:2021:biref ; Nakatsuka:2022 ; Naokawa:2023 ; Yin:2023:biref ; Naokawa:2024xhn ; Murai:2024yul . Measuring the spectral shape of the $EB$ power spectrum could break the degeneracy between the miscalibrated polarization angle and cosmic birefringence signals. Additional constraints on the late-time evolution of ALPs could be obtained by observing the polarized Sunyaev-Zel’dovich effect Lee:2022:pSZ-biref ; Namikawa:2023:pSZ and galaxy polarization Carroll:1997:radio ; Yin:2024:galaxy . These tomographic signals can help reduce degeneracies with the miscalibrated polarization angle.
@@
Anisotropies in cosmic birefringence are another critical observable that can confirm the isotropic signal. If the ALP field exhibits spatial fluctuations, $\delta\phi$, the resulting polarization rotation will be anisotropic (see, e.g., Carroll:1998:DE ; Lue:1999:biref-EB ; Caldwell:2011 ; Lee:2015 ; Leon:2017 ; Yin:2023:biref ; Ferreira:2023 ). Several studies have placed upper limits on these anisotropies by reconstructing the rotation angle Gluscevic:2012 ; PB15:rot ; BICEP2:2017lpa ; Contreras:2017 ; Namikawa:2020:biref ; SPT:2020:biref ; Gruppuso:2020 ; Bortolami:2022whx ; BK-LoS:2023 ; Zagatti:2024 , while others have constrained them using CMB polarization power spectra Li:2014 ; Alighieri:2014yoa ; Liu:2016dcg ; Zhang:2024dmi . However, time evolution tends to significantly suppress the $B$-mode power spectrum Namikawa:2024:BB .
@@
This paper is organized as follows. In Sec. II, we explain our reconstruction methodology. In Sec. III, we describe our data and simulation and test the estimator of anisotropic cosmic birefringence generated during reionization. Section IV shows our results for the reconstructed spectrum and the resulting constraint on the scale-invariant birefringence spectrum. We summarize our results and provide a forecast in Sec. V.
@@ II Tomographic reconstruction of anisotropic cosmic birefrigence
For a single source plane, such as the last-scattering surface of CMB, the rotation angle at each line-of-sight direction $\hat{\bm{n}}$ can be reconstructed from the off-diagonal mode-mode covariance between the $E$ and $B$ modes Kamionkowski:2009:derot ; Namikawa:2020:biref . The power spectrum of the anisotropic rotation angle can be obtained by squaring the rotation estimator and subtracting relevant biases Namikawa:2016:biref-est . Here, we extend the formalism to the case in which the CMB photons are generated at multiple epochs.
@@ II.1 Anisotropic cosmic birefringence from recombination and reionization
As a simple case, we first consider the case in which the CMB photons are generated from the recombination and reionization epochs. The presence of the cosmic birefringence effect rotates the primordial Stokes parameters as Greco:2022:cross ; Greco:2022:aniso-biref-tomography
@@
Here, $Q_{x}$ and $U_{x}$ represent the CMB polarization generated at reionization ($x={\rm rei}$) and recombination ($x={\rm rec}$). Similarly, $\alpha_{x}$ denotes the rotation angle at reionization ($x={\rm rei}$) and recombination ($x={\rm rec}$). Consequently, the rotation angle modifies the CMB $E$ and $B$ modes. The $E$ and $B$ modes are obtained by transforming the $Q$ and $U$ maps with the spin-2 spherical harmonics, $Y_{\ell m}^{\pm 2}$, as Kamionkowski:1996:eb ; Zaldarriaga:1996xe :
@@
Thus, the $E$ and $B$ modes in the presence of an anisotropic rotation angle are derived by substituting Eq. 2 into the above equation and are given up to linear order in $\alpha$ by
@@
The last term in the above denotes the Wigner $3j$-symbol. The off-diagonal elements of the covariance induced by the anisotropies of the rotation angle are given by
@@
Here, $\overline{E}_{\ell m}$ and $\overline{B}_{\ell m}$ are the observed multipoles filtered by their inverse variance. For a diagonal filtering, we obtain $\overline{X}_{\ell m}=\widehat{X}_{\ell m}/\widehat{C}^{\rm XX}_{\ell}$, where $X$ is either $E$ or $B$ and $\widehat{C}^{\rm XX}_{\ell}$ is the power spectrum of the observed multipoles, $\widehat{X}_{\ell m}$. In practice, however, the diagonal filtering is not optimal. In this paper, we use the optimal filtering as detailed in Eq. (2.11) of Ref. Lonappan:LiteBIRD-lens . Finally, we correct for the mean-field bias, $\langle\bar{\alpha}^{\rm rei}_{LM}\rangle$, and normalize to obtain the rotation angle:
@@
We can compute Eqs. 8 and 10 with a computationally efficient method as described in Appendix A of Namikawa:2020:biref . The entire code for reconstructing cosmic birefringence on the full sky is available at https://github.com/toshiyan/cmblensplus/tree/master.
@@ II.2 Anisotropic cosmic birefringence from multiple redshifts
Analogous to the estimator for $\alpha^{\rm rei}$, we define the estimator for $\alpha^{z}$ using Eq. (12) as follows:
@@
Here, $A_{L}^{z}$ is the same as the normalization defined in Eq. (10) but with $f^{\alpha,z}$ instead of $f^{\alpha,{\rm rei}}$. However, this estimator is biased because its expectation value includes contributions from anisotropic birefringence at different redshifts:
@@ III Data and simulations
Figure 1: Angular power spectrum of anisotropic cosmic birefringence at reionization measured from the Planck PR4 simulation with the $70\%$ Galactic mask, shown in blue. The CMB polarization and noise are derived from the Commander component separation method. The error bars denote the standard deviation with $400$ realizations. We also show the mean-field bias (orange) and disconnected bias, $N^{(0)}$ (green), which corresponds to the noise power spectrum of $\alpha^{\rm rei}_{LM}$.
Figure 2: Correlation coefficients of the angular power spectrum of anisotropic cosmic birefringence measured from the Planck PR4 simulation, as shown in Fig. 1. The lower-left triangle displays the results when the realization-dependent disconnected bias defined in Eq.(27) of Ref. Namikawa:2016:biref-est is used for estimating the disconnected bias, while the upper-left triangle shows, for comparison, the case where the simulation average of the disconnected bias defined in Eq.(26) of Ref. Namikawa:2016:biref-est is used.
@@
We use the Planck PR4 polarization data to reconstruct anisotropic cosmic birefringence, employing the curved-sky quadratic estimator defined in Eq. (9) to extract large-scale birefringence anisotropies.
The Planck PR4 polarization map is produced using two different component separation algorithms: Commander and SEVEM. We use both maps and the corresponding simulation sets to reconstruct anisotropic cosmic birefringence. Additionally, we generate a simulation where the CMB signal components are replaced with a signal that includes nonzero anisotropic cosmic birefringence, whose power spectrum follows a scale-invariant model. In this simulation, we first generate $\alpha_{LM}^{\rm rei}$ from a scale-invariant spectrum $C_{L}^{\alpha\alpha}=A_{\rm CB}^{\rm rei}\times 2\pi/L(L+1)$, and then rotate the CMB polarization signals at $\ell\leq 10$ with this rotation field. We assume $A_{\rm CB}^{\rm rei}=2\times 10^{-3}$.
@@
We evaluate the mean-field bias by averaging over the standard simulation realizations. Note that mean-field bias is also induced by the miscalibrated polarization angle Namikawa:2020:biref , which is not included in the simulations provided by the Planck Collaboration. We test this effect by rotating the polarization map in the simulation as we discuss in the next section.
In practice, partial sky coverage introduces mixing between different multipoles, and the diagonal normalization does not provide an accurate normalization. To correct for this, we follow the approach used in CMB lensing analyses. We first reconstruct $\alpha$ from a masked simulation that includes nonzero anisotropic birefringence. Next, we cross-correlate the input $\alpha$ with the reconstructed $\alpha$ and rescale the normalization at each $L$ so that the cross-power spectrum agrees with the input birefringence power spectrum.
From the reconstructed $\alpha$, we estimate the cosmic birefringence spectrum following Ref. Namikawa:2016:biref-est . The power spectrum of the estimator defined in Eq. 9 is a four-point function and is decomposed into the disconnected and connected parts. The disconnected part, $N_{L}^{(0)}$, is nonzero even in the absence of birefringence. We evaluate the disconnected bias using the realization-dependent approach as defined in Eq. (27) of Ref. Namikawa:2016:biref-est and subtract it from the power spectrum of the estimator. The N1 bias related to the connected part of the trispectrum is estimated using the simulations with nonzero birefringence. We also consider the residual power spectrum in the PR4 standard simulation as additional bias from the connected four-point function, $N^{\rm MC}_{L}$, arising from lensing Namikawa:2016:biref-est , partial-sky coverage, and noise. We subtract all these terms from the power spectrum of the estimator to extract the power spectrum of the anisotropic cosmic birefringence signal as described in Eq. (33) of Ref. Namikawa:2016:biref-est .
Figure 1 shows the angular power spectrum of anisotropic cosmic birefringence from the PR4 simulation, which contains no birefringence signals. We use the $70\%$ Galactic mask and simulation set with the Commander component separation method. We do not apply any normalization correction to the estimator. The estimated power spectrum is consistent with a null result. Figure 2 plots the correlation coefficients of the angular power spectrum. The realization-dependent method for $N_{L}^{(0)}$ successfully reduces the off-diagonal covariance. Figure 3 shows the angular power spectrum obtained from a simulation that includes anisotropic cosmic birefringence signals. We employ the $70\%$ Galactic mask and simulation set with the Commander component separation method. The power spectrum estimator accurately reproduces the input angular power spectrum. The mean-field, $N^{(1)}$, and $N^{\rm MC}$ are smaller than or comparable to the input signal with $A_{\rm CB}^{\rm rei}=2\times 10^{-3}$ which is much lower than the value constrained from the real data as we show in the next section. Note that $N^{(0)}$ is much larger than the other spectra and becomes approximately $100$ [deg2] at $L<10$ and $\mathcal{O}(10)$ [deg2] at $10\leq L\leq 100$. This large value of $N^{(0)}$ arises from the normalization correction applied to the estimator. It reflects a significant suppression of reconstructed anisotropic birefringence signals on large angular scales.